A motorist travels 160km at 80km/h and 160km at 100km/h. What is the average speed of the motorist for this trip?
1
Expert's answer
2020-11-02T09:24:17-0500

We can find the average speed of the motorist for this trip from the formula:


"v_{avg}=\dfrac{Total\ distance}{Total\ time\ elapsed}."


The total distance traveled is the sum of two distances: the distance "s_1" traveled at a speed 80 km/h  and the distance "s_2" traveled at a speed 100 km/h. Therefore, we can write:


"Total\ distance=s_1+s_2=160\ km+160\ km=320\ km."


The total time elapsed is the sum of two times: the time "t_1"that the motorist takes to travel the distance "s_1"and the time "t_2" that the motorist takes to travel the distance "s_2".

Therefore, we can write:


"Total\ time\ elapsed=t_1+t_2=\dfrac{s_1}{v_1}+\dfrac{s_2}{v_2},""Total\ time\ elapsed=\dfrac{160\ km}{80\ \dfrac{km}{h}}+\dfrac{160\ km}{100\ \dfrac{km}{h}}=3.6\ h."


Finally, we can find the average speed of the motorist for this trip:


"v_{avg}=\dfrac{Total\ distance}{Total\ time\ elapsed}=\dfrac{320\ km}{3.6\ h}=89 \ \dfrac{km}{h}."

Answer:

"v_{avg}=89 \ \dfrac{km}{h}."